OpenStudy (anonymous):

Find the largest (most positive) value of t satisfying: t2 + 16 t - 27 = 12 t - 6

OpenStudy (lgbasallote):

solve for t first. subtract 12t from both sides..what do you get?

OpenStudy (anonymous):

t^2+4t-27

OpenStudy (lgbasallote):

t^2 + 4t - 27 = -6 don't forget that right side now add 6 to both sides

OpenStudy (anonymous):

t^2+4t-21=0

OpenStudy (lgbasallote):

right. now factor out that equation

OpenStudy (anonymous):

how do you factor... thats what all my equations come to and i just dont remember how to do it

OpenStudy (lgbasallote):

if you're not used to factoring...you can just use the quadratic formula..do you know the formula?

OpenStudy (anonymous):

yes i found it let me try it out

OpenStudy (anonymous):

i got 4 ?

OpenStudy (lgbasallote):

no it's not 4

OpenStudy (lgbasallote):

let me give you a hint t^2 + 4t - 21 = 0 factor it out... (t+7)(t-3) = 0

OpenStudy (lgbasallote):

solve for t

OpenStudy (anonymous):

3 and -7

OpenStudy (lgbasallote):

so which one is the positive?

OpenStudy (anonymous):

3 Awesome so for this equation 3x^2-16x+16=0 it would be 16 and -1 because 16 mult by -1 is 16 so then it would be (x-1)(x+16)

OpenStudy (anonymous):

No, (x-1)(x+16) = x^2 +15x -16. You are looking for 3x^2-16x+16

OpenStudy (lgbasallote):

don't disregard that 3 beside x^2

OpenStudy (anonymous):

can i multiply the 16's by 3?

OpenStudy (lgbasallote):

just break down the middle term 3x^2 - 12x - 4x + 16 = 0 factor that out

OpenStudy (lgbasallote):

i have to go now so i hope you get this @kaylakelly13

OpenStudy (anonymous):

thank you so much

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

In FOIL: First Outside Inside Last You want the Last terms to equal 16 when multiplied. So, maybe it is (3x + 2)(x + 8). Because 2 x 8 = 16

OpenStudy (anonymous):

Any ideas?

OpenStudy (anonymous):

well wouldnt at least one need to be negative ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

But, 16 is positive. So what does that mean?

OpenStudy (anonymous):

that they are both negative

OpenStudy (anonymous):

Correct! So far you know this... (3x - )(x - ), fill in the blanks.

OpenStudy (anonymous):

I suggested 2 and 8 before: (3x - 2)(x - 8)

OpenStudy (anonymous):

could it be (3x-4)(x-4)?

OpenStudy (anonymous):

Multiply it out and tell me what you get.

OpenStudy (anonymous):

i got 3x^2-12x-4x+16 but i would get the same for (3x-2)(x-8)

OpenStudy (anonymous):

Let's look at (3x-2)(x-8): Use FOIL FIRST: 3x * x = 3x^2, that works OUTSIDE: 3x * -8 = -24x INSIDE: -2 * x = -2x LAST: -2 * -8 = 16 So, we get 3x^2 -24x -2x +16

OpenStudy (anonymous):

The OUTSIDE and INSIDE are not correct. So it looks like you were right: the answer is -4 and -4.

OpenStudy (anonymous):

but my answer has to be positive

OpenStudy (anonymous):

What do you mean? Why does it have to be positive?

OpenStudy (anonymous):

This is the original equation: 3x^2-16x+16=0

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

3x^2-16x+16 = (3x^2 - 4)(x - 4) = 3x^2-12x-4x+16 = 3x^2-16x+16

OpenStudy (anonymous):

Negative 4 times negative 4 = positive 16; is that what you mean?

OpenStudy (anonymous):

Or do you mean that x has to be positive?

OpenStudy (anonymous):

Find the largest (most positive) value of x satisfying

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

So, we know that 3x^2-16x+16 = (3x^2 - 4)(x - 4) right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

And (3x^2 - 4)(x - 4) = 0 in the original question

OpenStudy (anonymous):

So basically, for what values of x does (3x^2 - 4)(x - 4) = 0?

OpenStudy (anonymous):

Whenever you multiply a number by zero you get 0

OpenStudy (anonymous):

So, when (3x^2 - 4)(x - 4) = 0 (3x^2 - 4) could be zero (x - 4) could be zero

OpenStudy (anonymous):

Set (3x^2 - 4) equal to zero and solve for x Then set (x-4) equal to zero and solve for x The more positive x is your answer

OpenStudy (anonymous):

x - 4 = 0 Add 4 to both sides x = 4

OpenStudy (anonymous):

4 would be the answer cus the other one would be teh square root of 3/4

OpenStudy (anonymous):

3x^2 - 4 = 0 Add 4 to both sides 3x^2 = 4 Divide by 3 x^2 = 4/3 Square root x = SQRT(4/3), which simplifies into 2SQRT(3)/3 (DON't WORRY ABOUT THIS, IT'S LESS THAN 4).... So, x = 4 is largest

OpenStudy (anonymous):

thank you sooo much! this really helped me